# AED: Adaptable Error Detection for Few-shot Imitation Policy

### v1

@@ E.4 Pilot Study on Eeror Correction
$80.56\pm 11.65\%$ $80.56\pm 11.65\%$ $80.28\pm 11.60\%$ $71.67\pm 20.75\%$ $\mathbf{82.22\pm 10.17\%}$
We conducted the study on the Press Button task, which will most likely be corrected after an error occurs. In this task, the correction policy is defined as moving to the top center of the table. We let the DCT policy cooperate with the correction policy and four AED methods (SVDDED excluded), as summarized in Table 7. We have two findings from the results: (1) Our PrObe is verified to be the most accurate method again. In contrast, other AED methods will cause errors at the wrong timing, so the correction policy cannot improve performance (it may even have a negative impact on original successful trials). (2) The performance gain from the correction policy is minor since it lacks precise guidance in unseen environments.
We believe error correction in novel environments deserves a separate study due to its challenging nature, as observed in the pilot study. One potential solution is the human-in-the-loop correction, which operates through instructions (Sharma et al., 2022; Cui et al., 2023) or physical guidance (Li et al., 2021). However, their generalization ability and cost when applying to our AED task need further discussion and verification. We will leave this as a topic for our future work.
@@ Appendix F Limitations
Offline evaluation for the main experiment?
@@
No state-of-the-art (SOTA) baselines for the AED task?
As we formulate the novel AED task, there is no existing method that fits the task. We have found that some SOTA works (Li et al., 2021; Sharma et al., 2022; Cui et al., 2023; Liu et al., 2023a) also address the error detection or correction task in robotics. However, ours and theirs have different natures in terms of the task scenario. They usually require human-in-the-loop guidance, while in our setting, the robot is in an unseen environment, and the expert may only have the ability to perform demonstrations but is unable to correct the robot’s failures.
Additionally, we have made efforts to design baselines with different characteristics, as listed in Table 3. Our baselines include a traditional recurrent network, a one-class SVM, a metric-learning method, and a powerful transformer model. Notably, PrObe demonstrates superior performance compared to the baselines in our extensive experiments.
@@
Due to the practical challenges inherent in FSI tasks and the AED task, none of the existing robot manipulation benchmarks align with our specific requirements. For instance, FSI tasks necessitate multi-stage scenarios or variations in configurations and appearances between experts and agents. In the case of AED, it is crucial to gather failed rollouts in the base environments efficiently. As a result, we have had to develop our own environments and tasks. Moreover, the lack of real-world experiments is attributed to the following reasons: (1) Deploying the necessary resources for the AED task in the real world is high-cost and requires a long-term plan due to safety concerns. (2) The FSI policies still struggle to perform complex tasks in the real world, making assessments less than comprehensive. (3) Although PrObe achieves the best performance, there is still significant room for improvement in addressing the AED task.
@@
We emphasize that rashly conducting AED and FSI research in real environments may cause irreparable damage. Even previous error correction work (Wong et al., 2021) or related safe exploration research (Sootla et al., 2022; Anderson et al., 2023; Huang et al., 2023) was conducted in simulation environments, demonstrating that it is reasonable and valuable for us to conduct detailed verification in simulations first. As mentioned earlier, we have developed seven challenging and realistic multi-stage FSI tasks, each containing dozens of base and novel environments. To our knowledge, the simulation environments we established are the closest to the real scenarios in the literature.
Table 8: AED methods’ performance on our seven challenging FSI tasks. The value under the policy indicates the policy’s success rate in that task. Our proposed PrObe demonstrates its superiority among all the methods on both metrics and is the most robust method for different policies.
Metric [$\uparrow$] AED Close Drawer Press Button
NaiveDC DCT SCAN NaiveDC DCT SCAN
91.11% 80.56% 88.33% 51.94% 80.56% 75.56%
AUROC SVDDED 0.7404 0.8378 0.4079 0.4113 0.5710 0.4280
TaskEmbED 0.8395 0.7081 0.6498 0.4872 0.5429 0.4754
LSTMED 0.8186 0.6590 0.7867 0.3769 0.4886 0.4066
DCTED 0.8250 0.6413 0.7218 0.7194 0.7306 0.7491
PrObe 0.9133 0.7680 0.6978 0.7957 0.7506 0.7782
AUPRC SVDDED 0.2626 0.7039 0.1411 0.6956 0.3789 0.4653
TaskEmbED 0.6840 0.5995 0.2220 0.7280 0.3597 0.4657
LSTMED 0.7700 0.6493 0.6677 0.7051 0.4468 0.4487
DCTED 0.7813 0.6739 0.6390 0.8405 0.5734 0.6757
PrObe 0.8350 0.7438 0.5829 0.8851 0.7474 0.7505
Metric [$\uparrow$] AED Pick & Place Move Glass Cup
NaiveDC DCT SCAN NaiveDC DCT SCAN
55.00% 64.05% 71.19% 39.00% 51.80% 72.60%
AUROC SVDDED 0.7074 0.6125 0.5646 0.4447 0.6201 0.6473
TaskEmbED 0.6810 0.7523 0.6381 0.8370 0.5779 0.7950
LSTMED 0.7116 0.6555 0.7050 0.8086 0.6670 0.6804
DCTED 0.7493 0.7743 0.7482 0.7143 0.5954 0.6401
PrObe 0.7635 0.8173 0.8012 0.7487 0.6697 0.7425
AUPRC SVDDED 0.7971 0.7001 0.5078 0.8323 0.7188 0.6251
TaskEmbED 0.8028 0.8173 0.6065 0.9609 0.6823 0.8204
LSTMED 0.8148 0.7304 0.7103 0.9420 0.7720 0.6404
DCTED 0.8041 0.8294 0.7063 0.9012 0.7086 0.5713
PrObe 0.8665 0.8780 0.8072 0.9228 0.7907 0.7495
Metric [$\uparrow$] AED Organize Table Back to Box
NaiveDC DCT SCAN NaiveDC DCT SCAN
12.20% 79.00% 66.60% 08.89% 29.17% 58.89%
AUROC SVDDED 0.2581 0.5622 0.5000 0.4621 0.5537 0.5498
TaskEmbED 0.8078 0.6382 0.6193 0.3221 0.7220 0.7041
LSTMED 0.5946 0.7001 0.6734 0.4148 0.7707 0.8237
DCTED 0.5844 0.6702 0.5962 0.6022 0.4772 0.7489
PrObe 0.6808 0.6550 0.6759 0.8446 0.7276 0.8614
AUPRC SVDDED 0.8911 0.4800 0.5086 0.9367 0.8652 0.6433
TaskEmbED 0.9786 0.3922 0.4865 0.9375 0.9314 0.7401
LSTMED 0.9611 0.6352 0.6356 0.9555 0.9546 0.8913
DCTED 0.9452 0.5515 0.5255 0.9679 0.8534 0.7880
PrObe 0.9652 0.5905 0.6983 0.9903 0.9438 0.9097
Metric [$\uparrow$] AED Factory Packing Statistics
NaiveDC DCT SCAN Top 1 Avg.
45.42% 88.75% 63.75% counts Rank
AUROC SVDDED 0.3338 0.3849 0.6151 1 4.2
TaskEmbED 0.5564 0.7201 0.6916 3 3.2
LSTMED 0.6471 0.7934 0.7836 4 2.8
DCTED 0.5361 0.7509 0.5006 0 3.2
PrObe 0.6635 0.7670 0.8256 13 1.5
AUPRC SVDDED 0.5583 0.1600 0.5657 0 4.5
TaskEmbED 0.7057 0.5002 0.7287 3 3.5
LSTMED 0.7667 0.6335 0.8063 3 2.5
DCTED 0.6772 0.6224 0.6306 0 3.1
PrObe 0.7676 0.6759 0.8622 15 1.4
@@
Figure 15: AED methods’s ROC curves of all FSI tasks. The AUROC scores reported in Table 8 are the area under the curves.

### v2

@@ E.4 Pilot Study on Error Correction
[…]
To further examine the practicality of our AED task, we conducted a pilot study on error correction. In this experiment, the FSI policy is paused after the AED methods detect an error. Then, a correction policy from [19], which resets the agent to a predefined safe pose, is applied. Finally, the FSI policy continues to complete the task. We conducted the study on the Press Button task, where errors are most likely to occur and be corrected. Besides, the correction policy is defined as moving to the top center of the table. We allowed the DCT policy to cooperate with the correction policy and four AED methods (SVDDED excluded), as summarized in Table 8.
We have two following findings: (1) Our PrObe is verified to be the most accurate method once again. In contrast, other AED methods may cause errors at the wrong timing, making it challenging for the correction policy to improve performance (it may even have a negative impact on original successful trials). (2) The performance gain from the correction policy is minor, as it lacks precise guidance in unseen environments.
We believe that error correction in novel environments warrants a separate study due to its challenging nature, as observed in the pilot study. One potential solution is the human-in-the-loop correction, which operates through instructions [53, 54] or physical guidance [55]. However, their generalization ability and cost when applying to our AED task need further discussion and verification. We will leave this as a topic for our future work.
@@ Appendix F Limitations
Offline evaluation in AED inference?
@@
No comparison with state-of-the-art methods?
While there are no available state-of-the-art (SOTA) methods for the novel AED task we formulate, we have endeavored to design several strong baselines with various characteristics (cf. Section C of the Appendix). These baselines draw inspiration from relevant areas such as one-class classification, metric-learning, temporal modeling. Moreover, their number of learnable parameters is ten times that of our PrObe. Nonetheless, PrObe demonstrates superior performance compared to them in our extensive experiments.
Additionally, we observed a parallel line of research focused on detecting or correcting robot errors, with some SOTA approaches [55, 53, 54, 56]. However, our approach and theirs have different natures in terms of the task scenario. Their methods typically rely on human-in-the-loop guidance, whereas in our setting, the robot operates in an unseen environment, and the expert may only be able to provide demonstrations without the capability to correct the robot’s failures. Consequently, they are not suitable as baselines in this work.
@@
Due to the practical challenges inherent in FSI tasks and the AED task, we have had to develop our own benchmarks and tasks. Moreover, the lack of real-world experiments is attributed to the following reasons: (1) Deploying the necessary resources for the AED task in the real world is high-cost and requires a long-term plan due to safety concerns. (2) The FSI policies still struggle to perform complex tasks in the real world, making assessments less than comprehensive. (3) Although PrObe achieves the best performance, there is still significant room for improvement in addressing the AED task. We emphasize that hastily conducting AED and FSI research in real environments may cause irreparable damage.
@@
Even previous error correction work [19] or related safe exploration research [57, 58, 59] was conducted in simulation environments, demonstrating that it is reasonable and valuable for us to conduct detailed verification in simulations first. As mentioned earlier, we have developed seven challenging and realistic multi-stage FSI tasks, each containing dozens of base and novel environments. To our knowledge, the simulation environments we established are the closest to real-world scenarios in the literature.
Broader Impact
Our work focuses on addressing a critical learning task, specifically the error detection of policy behaviors, with the goal of accelerating the development of safety-aware robotic applications in real-world scenarios. While we have not identified any immediate negative impacts or ethical concerns, it is essential for us to remain vigilant and continuously assess potential societal implications as our research progresses.
@@
Figure 15: AED methods’ ROC curves for all FSI tasks. The AUROC scores reported in the main paper’s Figure 4 are the area under the curves.